What is the PV of a future sum if: FV = 15000 # of periods = 12 Interest rate = 5% Compounding times per period = 12
Answer #1

Ans $ 14269.92

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV = PV (1 + r )n
15000 = PV*(1+5%/12)^12
15000 = 1.05116189788173 * PV
PV = 15000 / 1.05116189788173
PV = 14269.92